# On the singular loci of higher secant varieties of Veronese embeddings

### v2

@@ 2.1. Projection from the incidence to the secant variety
Under the embedding $I\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\hookrightarrow{\mathbb{P}}^{l_{1}}$, taking $R\cap I$ in ${\mathbb{P}}^{l_{1}}$, we have $\operatorname{codim}(R\cap I,I)\geq 2$, as follows. Let $R^{\prime}$ be the center of the projection ${\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$, where $R^{\prime}\cap L=\emptyset$ and $\operatorname{codim}(R^{\prime},{\mathbb{P}}^{n})=m+1$. For $R^{\prime}(i)=\left\{(x_{1}^{\prime},\dots,x_{k}^{\prime})\in({\mathbb{P}}^{n})^{k}\mid x_{i}^{\prime}\in R^{\prime}\right\}$, we have
$R\cap I=q^{-1}\bigl{(}\bigcup_{i=1}^{k}R^{\prime}(i)\bigr{)},$
@@
where $q:I\rightarrow({\mathbb{P}}^{n})^{k}$. In particular, since dimension of fibers of $q|_{I^{0}}$ is constant, it follows that $\operatorname{codim}(R\cap I^{0},I)=m+1$. Next, let $\left\{W_{j}\right\}$ be the irreducible components of $I\setminus I^{0}$, and let $(a_{j},x_{j,1}^{\prime},\dots,x_{j,k}^{\prime})\in W_{j}$. A general $(n-m-1)$-plane ${\mathbb{P}}^{n-m-1}\subset{\mathbb{P}}^{n}$ satisfies $x_{j,i}^{\prime}\notin{\mathbb{P}}^{n-m-1}$. Since $\left\{W_{j}\right\}$ does not depend on the choice of $R^{\prime}={\mathbb{P}}^{n-m-1}$, we may assume $x_{j,i}^{\prime}\notin R^{\prime}$. Then $W_{j}\not\subset R$ for any $j$. Taking an irreducible component $Q$ of $R\cap I$ such that $Q\subset I\setminus I^{0}$, we have some $j$ such that $Q\subset R\cap W_{j}\subsetneq W_{j}\subsetneq I$. Hence $\operatorname{codim}(Q,I)\geq 2$. As a result, $\operatorname{codim}(R\cap I,I)\geq 2$.
@@
Now, let $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in I\setminus I^{0}$ satisfy $a\in{\mathbb{P}}^{\beta}$, where $\dim\langle\,x_{1},\dots,x_{k}\,\rangle<k-1$ for $x_{i}=v_{d}(x_{i}^{\prime})$ (note that we do not know $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$). We regard $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})$ as a point of ${\mathbb{P}}^{l_{1}}$ under the embedding $I\subset{\mathbb{P}}^{l_{1}}$.
@@
If $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\notin R$, then $\rho_{1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime})=(a,y_{1}^{\prime},\dots,y_{1}^{\prime})$ is determined in $\overline{\rho_{1}(I)}\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}$, where $y_{i}^{\prime}\in{\mathbb{P}}^{m}$ is the image of $x_{i}^{\prime}$ under ${\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{m}$. Since $a\in{\mathbb{P}}^{\beta}$, $\rho_{2}(\rho_{1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime}))=(a,y_{1}^{\prime},\dots,y_{1}^{\prime})$ is determined and is contained in $I_{(m)}$. Then, $\dim\langle\,y_{1},\dots,y_{k}\,\rangle<k-1$ for $y_{i}=v_{d}(y_{i}^{\prime})\in{\mathbb{P}}^{\beta}$, which means $(a,y_{1}^{\prime},\dots,y_{1}^{\prime})\in I_{(m)}\setminus I_{(m)}^{0}$.
@@
Assume $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in R$. Let ~$\tilde{{\mathbb{P}}}^{l_{1}}\subset{\mathbb{P}}^{l_{1}}\times{\mathbb{P}}^{l_{2}}$ be the graph of ${\mathbb{P}}^{l_{1}}\dashrightarrow{\mathbb{P}}^{l_{2}}$. In other words, $\widetilde{{{\mathbb{P}}}^{l_{1}}}=Bl_{R}({\mathbb{P}}^{l_{1}})$, the blowing-up of ${\mathbb{P}}^{l_{1}}$ with respect to $R$. Let $r_{1}:\tilde{{\mathbb{P}}}^{l_{1}}\rightarrow{\mathbb{P}}^{l_{1}}$ and $r_{2}:\tilde{{\mathbb{P}}}^{l_{1}}\rightarrow{\mathbb{P}}^{l_{2}}$ be projections. Note that, for the strict transformation $S\subset\tilde{{\mathbb{P}}}^{l_{1}}$ of ${\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\subset{\mathbb{P}}^{l_{1}}$, two composite morphisms $S\xrightarrow{r_{1}|_{S}}{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{N}$ and $S\xrightarrow{r_{2}|_{S}}{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\rightarrow{\mathbb{P}}^{N}$ coincide since it holds on an open subset of $S$. Let $E\subset\tilde{{\mathbb{P}}}^{l_{1}}$ be the exceptional divisor, and let $\widetilde{I}\subset\tilde{{\mathbb{P}}}^{l_{1}}$ be the strict transformation of $I\subset{\mathbb{P}}^{l_{1}}$. Then $r_{2}(\widetilde{I})=\overline{\rho_{1}(I)}$.
@@
Let $E_{1}=r_{1}^{-1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\cap(E\cap\widetilde{I})$, the fiber of $E\cap\widetilde{I}\rightarrow R\cap I$ at $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})$. It follows from $\operatorname{codim}(R\cap I,I)\geq 2$ that $\dim(E_{1})\geq 1$. Since $r_{1}^{-1}(z)\simeq r_{2}(r_{1}^{-1}(z))\subset{\mathbb{P}}^{l_{2}}$ for each $z\in{\mathbb{P}}^{l_{1}}$, we have $\dim(r_{2}(E_{1}))\geq 1$. Since the image of $r_{2}(E_{1})$ under ${\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\rightarrow{\mathbb{P}}^{N}$ is $\left\{a\right\}$, there is $C\subset({\mathbb{P}}^{m})^{k}$ of positive dimension such that $r_{2}(E_{1})=\left\{a\right\}\times C\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}$. Since $a\in{\mathbb{P}}^{\beta}$ and $\rho_{2}(r_{2}(E_{1}))\subset\rho_{2}(r_{2}(\tilde{I}))=\overline{\rho_{2}(\rho_{1}(I))}=I_{(m)}$, we have $\left\{a\right\}\times C=\rho_{2}(\left\{a\right\}\times C)\subset I_{(m)}$. ∎
@@
Let $a\in\sigma_{k}(v_{d}(L))\cap\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (note that $a$ can be in the boundary of $\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$). For a general point $b_{0}\in\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, we take an irreducible curve $C\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ such that $a,b_{0}\in C$. Let $\pi_{2}:{\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$ be the linear projection in Remark 8 (a). Since $a\in{\mathbb{P}}^{\beta}$, we have $a=\pi_{2}(a)\in C^{\prime}:=\overline{\pi_{2}(C)}\subset{\mathbb{P}}^{\beta}$.
Since $b_{0}$ is general, for a general point $b\in C$, we have $b\in\langle\,x_{1},\dots,x_{k-1}\,\rangle$ with $x_{1},\dots,x_{k-1}\in v_{d}({\mathbb{P}}^{n})$. Take $x_{i}^{\prime}\in{\mathbb{P}}^{n}$ with $x_{i}=v_{d}(x_{i}^{\prime})$. As in Remark 8 (a), we take $y_{i}^{\prime}\in L$ for each $i=1,\dots,k-1$ and say $y_{i}=v_{d}(y_{i}^{\prime})$. Then $\pi_{2}(x_{i})=y_{i}$. It follows that $\pi_{2}(b)\in\langle\,y_{1},\dots,y_{k-1}\,\rangle\subset\sigma_{k-1}(v_{d}(L))$. As a result, $a\in C^{\prime}\subset\sigma_{k-1}(v_{d}(L))$ and the assertion follows. ∎
@@
    as a set. For one inclusion is obvious, let us prove $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\cap\langle v_{d}({\mathbb{P}}^{m})\rangle\subset\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$. Suppose that it does not hold. Then, there exists a form $f\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\cap\langle v_{d}({\mathbb{P}}^{m})\rangle$ with $f\in\sigma_{k_{0}}(v_{d}({\mathbb{P}}^{m}))\setminus\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ for some $k_{0}>k$. Then, $f\in\sigma_{k_{0}}(v_{d}({\mathbb{P}}^{m}))\cap\sigma_{k_{0}-1}(v_{d}({\mathbb{P}}^{n}))$ so that $f\in\sigma_{k_{0}-1}(v_{d}({\mathbb{P}}^{m}))$ by Lemma 10. Similarly, as repeating the same ‘descent’ argument, we have $f\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$, which is a contradiction. Thus, the equality in (4) is true.
    In other words, for any $d$-th Veronese embedding $X=v_{d}({\mathbb{P}}^{n})$ and the linear span $L=\langle v_{d}({\mathbb{P}}^{m})\rangle$ of any sub-Veronese variety $v_{d}({\mathbb{P}}^{m})$, we showed that $(X,L)$ is a border rank preserving pair for any $k,d>0$ in the terminology of [21, definition 5.7.3.1] (we’d also like to note that [5, theorem 1.1] can imply the same result for any $d\geq k$ in case of ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$).
@@
    (Every $\Sigma_{k,d}(m)$ is closed) Recall that the symmetric subspace variety $Sub_{m}(S^{d}V)$ ([21, section 7.1.3]) is defined as
@@ 2.2. General secant fiber of a subsecant variety, entry locus, and its Veronese image
(5) $J^{0}:=\left\{(a,x_{1},\dots,x_{k})\mid a\in\langle\,x_{1},\dots,x_{k}\,\rangle\text{ and }\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1\right\},$
with the projections $p:J\rightarrow\sigma_{k}(Z)\subset{\mathbb{P}}^{\beta}$ and
$q_{i}:J\rightarrow(Z)^{k}\rightarrow Z,$
@@
For $a\in\sigma_{k}(Z)$, the scheme-theoretic image $q_{i}(p^{-1}(a))$ in $Z$ is called the ($k$-th) entry locus of $Z$ with respect to $a$ in the literature. It is known that for a general $a\in\sigma_{k}(Z)$ the locus $q_{i}(p^{-1}(a))$ is equidimensional and moreover, if $Z$ is smooth and in the characteristic 0, then $p^{-1}(a)$ is smooth so that $q_{i}(p^{-1}(a))$ is reduced (see [28, definition 1.4.5]).
@@
$\mathbb{T}_{x}(X)\subset\mathbb{T}_{a}(\sigma_{k}(X))~{},$
@@
Since $\sigma_{k}(Z)\setminus\operatorname{Sing}(\sigma_{k}(X))$ is non-empty open in $\sigma_{k}(Z)$ and $a$ is general, it is a smooth point of $\sigma_{k}(X)$, and also is contained in the $(k-1)$-plane $\langle\,x_{1},\dots,x_{k}\,\rangle$ for general $x_{1},\dots,x_{k}\in Z$ with $x_{i}=x$. Then the assertion follows by Terracini’s lemma (cf. [28, corollary 1.4.2], [29, Chapter II, 1.10]). ∎
@@
For a projective variety $Z$, let $\langle\,Z\,\rangle={\mathbb{P}}^{\beta}$ and consider the incidence $J$ as (5). Suppose that the ($k-1$)-secant of $Z$ is not defective and not equal to ${\mathbb{P}}^{\beta}$. Then, for a general point $(a,x_{1},\dots,x_{k})\in J$, and for any irreducible component $F$ of $p^{-1}(a)$ containing $(a,x_{1},\dots,x_{k})$, $q_{i}|_{F}:F\rightarrow q_{i}(F)$ is generically finite.
@@
Let $J_{1}=q_{1}^{-1}(x_{1})\subset J$. For the projection $q_{2,\dots,k}:J\rightarrow(Z)^{k-1},(a,\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\mapsto(\tilde{x}_{2},\dots,\tilde{x}_{k})$, a general fiber of $q_{2,\dots,k}|_{J_{1}}:J_{1}\rightarrow(Z)^{k-1}$ at $(\tilde{x}_{2},\dots,\tilde{x}_{k})$ is isomorphic to the ($k-1$)-plane $\langle\,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$. In particular, $\pi_{x_{1}}\circ p|_{J_{1}}:J_{1}\rightarrow\pi_{x_{1}}(\sigma_{k-1}(Z))$ is surjective. Let $\alpha\in\pi_{x_{1}}(\sigma_{k-1}(Z))$ be a general point. Then we may consider the element $(a,x_{1},x_{2},\dots,x_{k})$ as a general point in $J_{1}$ with $\alpha=\pi_{x_{1}}(a)$.
Let $F$ be an irreducible component of $p^{-1}(a)$ containing $(a,x_{1},x_{2},\dots,x_{k})$, and suppose that $q_{1}|_{F}$ has fibers of positive dimensions. Then, the fiber of $q_{1}|_{F}$ at $x_{1}\in q_{1}(F)$,
@@
is of positive dimension. For $(a,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in F\cap J_{1}$, considering a line $\langle\,a,x_{1}\,\rangle\subset\langle\,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$ and a hyperplane $\langle\,\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$, we may take a point $\tilde{b}\in\langle\,a,x_{1}\,\rangle\cap\langle\,\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$. Note that $\pi_{x_{1}}(\tilde{b})=\pi_{x_{1}}(a)=\alpha$. It gives the following locus of positive dimension,
$\left\{(\tilde{b},\tilde{x}_{2},\dots,\tilde{x}_{k})\mid(a,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in F\cap J_{1}\right\}\subset\rho^{-1}(\alpha)\subset J^{\prime},$

### v3

@@ 2.1. Projection from the incidence to the secant variety
Under the embedding $I\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\hookrightarrow{\mathbb{P}}^{l_{1}}$, taking $R\cap I$ in ${\mathbb{P}}^{l_{1}}$, we have $\operatorname{codim}(R\cap I,I)\geq 2$, as follows. For the center $\hat{L}$ of $\pi_{1}:{\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$, we set
$\hat{L}(i)=\{\,(x_{1}^{\prime},\dots,x_{k}^{\prime})\in({\mathbb{P}}^{n})^{k}\mid x_{i}^{\prime}\in\hat{L}\,\}.$
We consider the projection $\bar{q}:{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\rightarrow({\mathbb{P}}^{n})^{k}$. Then a point $z\in{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}$ belongs to the indeterminacy locus of $\rho_{1}$ if and only if $\bar{q}(z)\in\hat{L}(i)$ for some $i$ with $1\leq i\leq k$. Since $\pi_{R}|_{I}(=\pi_{R}\circ Seg_{1}|_{I})$ coincides with $Seg_{2}\circ\rho_{1}|_{I}$, and since $R\cap I$ is the indeterminacy locus of $\pi_{R}|_{I}$, we have
$R\cap I=q^{-1}\bigl{(}\bigcup_{i=1}^{k}\hat{L}(i)\bigr{)},$
@@
where $q$ is equal to $\bar{q}|_{I}:I\rightarrow({\mathbb{P}}^{n})^{k}$. In particular, since the dimension of the fibers of $q|_{I^{0}}$ is constant, and since $\operatorname{codim}(\hat{L}(i),({\mathbb{P}}^{n})^{k})=m+1$, it follows that $\operatorname{codim}(R\cap I^{0},I)=m+1$. Now, let $\left\{Q_{s}\right\}$ be the irreducible components of $R\cap I$. Then $\operatorname{codim}(Q_{s},I)=m+1$ in the case $Q_{s}\cap I^{0}\neq\emptyset$, i.e., $Q_{s}\not\subset I\setminus I^{0}$. In order to consider the remaining case $Q_{s}\subset I\setminus I^{0}$, we use the irreducible components $\left\{W_{j}\right\}$ of $I\setminus I^{0}$. For $(a_{j},x_{j,1}^{\prime},\dots,x_{j,k}^{\prime})\in W_{j}$ given in (6), since $x_{j,i}^{\prime}\notin\hat{L}$, it follows $W_{j}\not\subset R$ for any $j$.
If an irreducible component $Q_{s}$ of $R\cap I$ satisfies $Q_{s}\subset I\setminus I^{0}$, then there is some $j$ such that $Q_{s}\subset R\cap W_{j}\subsetneq W_{j}\subsetneq I$, and then $\operatorname{codim}(Q_{s},I)\geq 2$. As a result, in any case, each irreducible component of $R\cap I$ is of codimension $m+1$ or at least $\geq 2$, which implies $\operatorname{codim}(R\cap I,I)\geq 2$.
@@
Now, let $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in I\setminus I^{0}$ satisfy $a\in{\mathbb{P}}^{\beta}$, where $\dim\langle\,x_{1},\dots,x_{k}\,\rangle<k-1$ for $x_{i}=v_{d}(x_{i}^{\prime})$ as in Remark 7 (a) (note that we do not know $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$). We regard $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})$ as a point of ${\mathbb{P}}^{l_{1}}$ under the embedding $I\subset{\mathbb{P}}^{l_{1}}$.
@@
If $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\notin R$, then $\rho_{1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime})=(a,y_{1}^{\prime},\dots,y_{1}^{\prime})$ is determined in $\overline{\rho_{1}(I)}\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}$, where $y_{i}^{\prime}\in{\mathbb{P}}^{m}$ is the image of $x_{i}^{\prime}$ under ${\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{m}$. Since $a\in{\mathbb{P}}^{\beta}$,
$\rho_{2}(\rho_{1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime}))=(a,y_{1}^{\prime},\dots,y_{1}^{\prime})$
is determined and is contained in $I_{(m)}$. Then, $\dim\langle\,y_{1},\dots,y_{k}\,\rangle<k-1$ for $y_{i}=v_{d}(y_{i}^{\prime})\in{\mathbb{P}}^{\beta}$, which means $(a,y_{1}^{\prime},\dots,y_{1}^{\prime})\in I_{(m)}\setminus I_{(m)}^{0}$.
@@
Assume $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in R$. We consider the blowing-up ~$\tilde{{\mathbb{P}}}^{l_{1}}$ of ${\mathbb{P}}^{l_{1}}$ with respect to $R$, and the projections $r_{1},r_{2}$ in the diagram (7). Note that, for the strict transformation $S\subset\tilde{{\mathbb{P}}}^{l_{1}}$ of ${\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\subset{\mathbb{P}}^{l_{1}}$, two composite morphisms
$S\xrightarrow{r_{1}|_{S}}{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{N}\,\text{ and }\,S\xrightarrow{r_{2}|_{S}}{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\rightarrow{\mathbb{P}}^{N}$
coincide since it holds on an open subset of $S$. Let $E\subset\tilde{{\mathbb{P}}}^{l_{1}}$ be the exceptional divisor, and let $\widetilde{I}\subset\tilde{{\mathbb{P}}}^{l_{1}}$ be the strict transformation of $I\subset{\mathbb{P}}^{l_{1}}$. Then $r_{2}(\widetilde{I})=\overline{\rho_{1}(I)}$. Let
$E_{1}=r_{1}^{-1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\cap(E\cap\widetilde{I}),$
@@
the fiber of $E\cap\widetilde{I}\rightarrow R\cap I$ at $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})$. It follows from $\operatorname{codim}(R\cap I,I)\geq 2$ that $\dim(E_{1})\geq 1$. Since $r_{1}^{-1}(z)\simeq r_{2}(r_{1}^{-1}(z))\subset{\mathbb{P}}^{l_{2}}$ for each $z\in{\mathbb{P}}^{l_{1}}$, we have $\dim(r_{2}(E_{1}))\geq 1$. Since the image of $r_{2}(E_{1})$ under ${\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\rightarrow{\mathbb{P}}^{N}$ is $\left\{a\right\}$, there is $C\subset({\mathbb{P}}^{m})^{k}$ of positive dimension such that $r_{2}(E_{1})=\left\{a\right\}\times C\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}$. Since $a\in{\mathbb{P}}^{\beta}$ and
$\rho_{2}(r_{2}(E_{1}))\subset\rho_{2}(r_{2}(\tilde{I}))=\overline{\rho_{2}(\rho_{1}(I))}=I_{(m)},$
we have $\left\{a\right\}\times C=\rho_{2}(\left\{a\right\}\times C)\subset I_{(m)}$. ∎
@@
Let $a\in\sigma_{k}(v_{d}(L))\cap\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (note that $a$ can be in the boundary of $\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$). For a general point $b_{0}\in\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, we take an irreducible curve $C\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ such that $a,b_{0}\in C$. Let $\pi_{2}:{\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$ be the linear projection in Remark 8 (a), and let $C^{\prime}=\overline{\pi_{2}(C)}\subset{\mathbb{P}}^{\beta}$. Since $a\in{\mathbb{P}}^{\beta}$, we have $a=\pi_{2}(a)\in C^{\prime}$.
Since $b_{0}$ is general, for a general point $b\in C$, we have $b\in\langle\,x_{1},\dots,x_{k-1}\,\rangle$ with $x_{1},\dots,x_{k-1}\in v_{d}({\mathbb{P}}^{n})$. Take $x_{i}^{\prime}\in{\mathbb{P}}^{n}$ with $x_{i}=v_{d}(x_{i}^{\prime})$. Setting $y_{i}=\pi_{2}(x_{i})$, we have $y_{i}=v_{d}(y_{i}^{\prime})$ with $y_{i}^{\prime}\in L$ for each $i=1,\dots,k-1$ as in Remark 8 (a). It follows that $\pi_{2}(b)\in\langle\,y_{1},\dots,y_{k-1}\,\rangle\subset\sigma_{k-1}(v_{d}(L))$. As a result, $a\in C^{\prime}\subset\sigma_{k-1}(v_{d}(L))$ and the assertion follows. ∎
@@
    as a set. Since one inclusion is obvious, let us prove $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\cap\langle v_{d}({\mathbb{P}}^{m})\rangle\subset\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$. Suppose that it does not hold. Then, there exists a form $f\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\cap\langle v_{d}({\mathbb{P}}^{m})\rangle$ with $f\in\sigma_{k_{0}}(v_{d}({\mathbb{P}}^{m}))\setminus\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ for some $k_{0}>k$. Then, $f\in\sigma_{k_{0}}(v_{d}({\mathbb{P}}^{m}))\cap\sigma_{k_{0}-1}(v_{d}({\mathbb{P}}^{n}))$ so that $f\in\sigma_{k_{0}-1}(v_{d}({\mathbb{P}}^{m}))$ by Lemma 10. Similarly, as repeating the same ‘descent’ argument, we have $f\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$, which is a contradiction. Thus, the equality in (8) is true.
    In other words, for any $d$-th Veronese embedding $X=v_{d}({\mathbb{P}}^{n})$ and the linear span $L=\langle v_{d}({\mathbb{P}}^{m})\rangle$ of any sub-Veronese variety $v_{d}({\mathbb{P}}^{m})$, we showed that $(X,L)$ is a border rank preserving pair for any $k,d>0$ in the terminology of [Land12, Definition 5.7.3.1] (we would also like to note that [BGL13, Theorem 1.1] can imply the same result for any $d\geq k$ in case of ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$).
@@
    (Every $\Sigma_{k,d}(m)$ is closed) Recall that the symmetric subspace variety $Sub_{m}(S^{d}V)$ ([Land12, section 7.1.3]) is defined as
@@ 2.2. General secant fiber of a subsecant variety, entry locus, and its Veronese image
(9) $J^{0}=\left\{(a,x_{1},\dots,x_{k})\mid a\in\langle\,x_{1},\dots,x_{k}\,\rangle\text{ and }\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1\right\},$
with the projections
$p:J\rightarrow\sigma_{k}(Z)\subset{\mathbb{P}}^{\beta},\quad q_{i}:J\rightarrow(Z)^{k}\rightarrow Z,$
@@
For $a\in\sigma_{k}(Z)$, the scheme-theoretic image $q_{i}(p^{-1}(a))$ in $Z$ is called the ($k$-th) entry locus of $Z$ with respect to $a$ in the literature. It is known that for a general $a\in\sigma_{k}(Z)$ the locus $q_{i}(p^{-1}(a))$ is equidimensional and moreover, if $Z$ is smooth and in the characteristic 0, then $p^{-1}(a)$ is smooth so that $q_{i}(p^{-1}(a))$ is reduced (see [Rus16, Definition 1.4.5]).
@@
$\mathbb{T}_{x}(X)\subset\mathbb{T}_{a}(\sigma_{k}(X)),$
@@
Since $\sigma_{k}(Z)\setminus\operatorname{Sing}(\sigma_{k}(X))$ is non-empty open in $\sigma_{k}(Z)$ and $a$ is general, it is a smooth point of $\sigma_{k}(X)$; hence the embedded tangent space $\mathbb{T}_{a}(\sigma_{k}(X))$ is defined. In addition, $a$ is contained in the $(k-1)$-plane $\langle\,x_{1},\dots,x_{k}\,\rangle$ for general $x_{1},\dots,x_{k}\in Z$ with $x_{i}=x$. Then the assertion follows by Terracini’s lemma (cf. [Rus16, Corollary 1.4.2], [Zak93, Chapter II, 1.10, Chapter V, 1.4]). ∎
@@
For a projective variety $Z$, let $\langle\,Z\,\rangle={\mathbb{P}}^{\beta}$ and consider the incidence $J$ as (9) with $k\geq 2$. Suppose that the ($k-1$)-secant of $Z$ is not defective and not equal to ${\mathbb{P}}^{\beta}$. Then, for a general point $(a,x_{1},\dots,x_{k})\in J$, and for any irreducible component $F$ of $p^{-1}(a)$ containing $(a,x_{1},\dots,x_{k})$, $q_{i}|_{F}:F\rightarrow q_{i}(F)$ is generically finite.
@@
Let $J_{1}=q_{1}^{-1}(x_{1})\subset J$. Then $\pi_{x_{1}}\circ p|_{J_{1}}:J_{1}\rightarrow\pi_{x_{1}}(\sigma_{k-1}(Z))$ is dominant (this is because, for general $b\in\sigma_{k-1}(Z)$, we take a general point $c\in\langle\,x_{1},b\,\rangle$ and take $k-1$ points $\tilde{x}_{2},\dots,\tilde{x}_{k}\in Z$ such that $b\in\langle\,\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$; then $(c,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in J_{1}$, whose image under $\pi_{x_{1}}\circ p$ is $\pi_{x_{1}}(c)=\pi_{x_{1}}(b)$). Since $(a,x_{1},\dots,x_{k})\in J_{1}$ is general in $J$ and by the dominance of $\pi_{x_{1}}\circ p|_{J_{1}}$, we may consider $\alpha=\pi_{x_{1}}(a)$ as a general point in $\pi_{x_{1}}(\sigma_{k-1}(Z))$.
Let $F$ be an irreducible component of $p^{-1}(a)$ containing $(a,x_{1},x_{2},\dots,x_{k})$, and suppose that $q_{1}|_{F}$ has general fibers of positive dimensions. Then, the fiber of $q_{1}|_{F}$ at $x_{1}\in q_{1}(F)$,
@@
is of positive dimension, which means that we have $(a,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in F\cap J_{1}$ for fixed $a,x_{1}$ and moving $(\tilde{x}_{2},\dots,\tilde{x}_{k})$ with positive dimension.
For general $(a,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in F\cap J_{1}$, we have the intersection point $b(\tilde{x}_{2},\dots,\tilde{x}_{k})$ of the line $\langle\,a,x_{1}\,\rangle$ and the hyperplane $\langle\,\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$ in $\langle\,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle={\mathbb{P}}^{k-1}$. Note that
$\pi_{x_{1}}(b(\tilde{x}_{2},\dots,\tilde{x}_{k}))=\pi_{x_{1}}(a)=\alpha.$
The $k$-tuple $(b(\tilde{x}_{2},\dots,\tilde{x}_{k}),\tilde{x}_{2},\dots,\tilde{x}_{k})$ moves with positive dimension since so does ($k-1$)-tuple $(\tilde{x}_{2},\dots,\tilde{x}_{k})$. In other words, the following locus is of positive dimension,
$\left\{(b(\tilde{x}_{2},\dots,\tilde{x}_{k}),\tilde{x}_{2},\dots,\tilde{x}_{k})\mid(a,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in F\cap J_{1}\right\}\subset\rho^{-1}(\alpha)\subset J^{\prime},$